# Denoising Diffusion Models for Inpainting of Healthy Brain Tissue

### v1

@@ Abstract
Our approach builds on denoising diffusion probabilistic models. We use a 2D model that is trained using slices in which healthy tissue was cropped out and is learned to be inpainted again. This allows us to use the ground truth healthy tissue during training. In the sampling stage, we replace the slices containing diseased tissue in the original 3D volume with the slices containing the healthy tissue inpainting. With our approach, we achieve comparable results to the competing methods. On the validation set our model achieves a mean SSIM of $0.7804$, a PSNR of $20.3525$ and a MSE of $0.0113$. In future we plan to extend our 2D model to a 3D model, allowing to inpaint the region of interest as a whole without losing context information of neighboring slices.
@@ 2.2 Inpainting of masked region
Figure 2: Overview of the training process. Context information is provided through the concatenation of input $b_{i}$ and mask $m_{i}$ with the noisy image $x_{i,t}$, which originated from the ground truth $x_{0}$ = $x_{i,GT}$ using Equation 2. The concatenated image $X_{t}$ is used by the diffusion model to predict a slightly denoised image $x_{i,t-1}$ from $x_{i,t}$ using Equation 5.
@@ 2.2.1 Training
where $\epsilon_{\theta}(X_{t},t)$ is the model output at time step $t$, $\sigma_{t}$ is the variance scheme and $\mathbf{z}$ covers the stochastic component of the process. As shown in [14], to train the diffusion model $\epsilon_{\theta}$ the following loss term with $\epsilon$ $\sim$ $\mathcal{N}(0,\mathbf{I})$ is used:
@@ 3.1 Data Preprocessing
The data set consists of T1 scans from the multi-modal BraTS 2022 glioma segmentation challenge [4, 5, 6, 9]. The data is a retrospective collection of brain tumor MR scans originating from multiple different institutions using different equipment and imaging protocols but all working under standard clinical conditions. For each patient, the challenge organizers provide T1 scans of brains affected by tumors, masks of the tumor regions, masks of healthy tissue regions, combined masks of the tumor and healthy regions as well as "T1-voided" scans, in which the combined masks are cropped out from the original T1 image. For our approach, we use the T1 images and the healthy region masks during training. As our final goal is to learn to replace healthy tissue, for training, we crop out the healthy masks from the T1 images to obtain a masked T1 image in which only a healthy region is missing. We remove the top and bottom 0.1 percentile of voxel intensities and normalize the scans to voxel values between 0 and 1. As our model is constructed for 2D data, we slice each 3D scan into transversal slices that are then cropped to a size of [224, 224]. Only background pixels are affected by the cropping. For training, we only use slices showing a non-zero mask. The provided training set consists of 1251 subjects which were used for a two-fold cross-validation, each using 1241 subjects for training and 10 for validation.
@@ 3.3 Postprocessing and Evaluation
After sampling, we apply a Gaussian filter with a standard deviation $sigma=1.075$ for the Gaussian kernel to smoothen borders between the slices. In addition to the training data, the challenge organizers provided a validation set of 219 subjects that could be evaluated using our model and submitted to the Synapse [17] platform to obtain structural similarity index measure (SSIM), peak signal-to-noise ratio (PSNR) and mean squared error (MSE). For submission to the Synapse platform, each model output is normalized to the intensity range of the corresponding input image from which the first and last 0.5 percentile were clipped. After submission of the final model as a MLCube [18], it will be evaluated on a non-public test set of several hundred cases.
@@ 4.1 Cross-validation using Train-Validation Splits from Training Data Set
Figure 4: Exemplary 2D transversal slices. The 2D generated samples are of good quality. Moreover, they show a high similarity to the corresponding ground truth.
Figure 4 provides two examples of a given 2D masked image, the corresponding mask, the generated sample and the ground truth in the transversal view. In 2D, our approach manages to generate consistent results. However, if we look at the whole 3D volume, stripe artifacts can appear in the sagittal and coronal plane if multiple sequential slices are replaced within the volume. This is shown in the bottom left part of Figure 5. One possibility to reduce the visibility of these artifacts is the application of a Gaussian filter, shown in the bottom right part of Figure 5.
Figure 5: Exemplary images of the validation set in the coronal (left) and sagittal (right) view. We observe stripe artifacts due to the stacking of multiple transversal slices. These artifacts can be reduced by applying a Gaussian filter.
@@ 4.2 Validation on Synapse Server
Table 2: Results of the evaluation of our model’s performance on the validation set. Evaluated on the Synapse server.
@@ 5 Discussion
We present an approach for healthy tissue inpainting based on 2D DDPMs. Taking advantage of the excellent image synthesis performance of DDPMs, we apply these models to masked brain MR images. Image context of the masked image is provided by concatenating the masked image to the input of the diffusion model in every time step $t$. These concatenated images are created to guide the denoising process. As only healthy tissue is cropped out, the ground truth is available during training. Regarding the validation, we achieve similar results in terms of SSIM, PSNR and MSE as the other challenge participants. Our model creates good results in 2D, but a drawback of our method is that the stacking of the 2D slices can lead to stripe artifacts, which we tried to reduce using a Gaussian filter. As an alternative, we tried to avoid the stripe artifacts using a sequential approach, where we included the previous ground truth slice as an additional channel that was concatenated with the noisy image during training. During sampling, the previously sampled slice was concatenated with the noisy image. However, this approach still needs modifications, as it seemed as if the errors accumulated over the sequence of slices. An advantage of the model we used for this challenge is that due to the two-dimensionality, it needs less computational resources than 3D models and shows a stable training progression. Nevertheless, in future, we aim to expand our approach to a 3D model, e.g., similar to [19]. So far we experimented with a 3D model using the same framework as presented in this paper, but it was limited regarding memory consumption, which led to limitations in network architecture tuning. A drawback of diffusion models in general is the slow sampling speed compared to other generative models. We aim to improve this using e.g., Heun sampling schemes [20] or consistency models [21]. We conclude that our method provides a baseline that shows a good performance for the task and has potential to be fine-tuned using different architectural and sampling changes.

### v2

@@ Abstract
Our approach builds on denoising diffusion probabilistic models. We use a 2D model that is trained using slices in which healthy tissue was cropped out and is learned to be inpainted again. This allows us to use the ground truth healthy tissue during training. In the sampling stage, we replace the slices containing diseased tissue in the original 3D volume with the slices containing the healthy tissue inpainting. With our approach, we achieve the second place in the challenge. On the test set our model achieves a mean SSIM of $0.8271$, a PSNR of $20.4949$ and a MSE of $0.0115$. In future we plan to extend our 2D model to a 3D model, allowing to inpaint the region of interest as a whole without losing context information of neighboring slices.
@@ 2.2 Inpainting of masked region
Figure 2: Overview of the training process. Context information is provided through the concatenation of input $b_{i}$ and mask $m_{i}$ with the noisy image $x_{i,t}$, which originated from the ground truth $x_{0}$ = $x_{i,GT}$ using Equation 2. The concatenated image $X_{t}$ is used by the diffusion model to predict a slightly denoised image $x_{i,t-1}$ from $x_{i,t}$ using Equation 5. The diffusion model is trained using an MSE loss between the real and the predicted noise, summarized in Equation 6.
@@ 2.2.1 Training
where $\epsilon_{\theta}(X_{t},t)$ is the diffusion model output at time step $t$, $\sigma_{t}$ is the variance scheme and $\mathbf{z}$ covers the stochastic component of the process. As shown in [14], to train the diffusion model $\epsilon_{\theta}$ the following loss term with $\epsilon$ $\sim$ $\mathcal{N}(0,\mathbf{I})$ is used:
@@ 3.1 Data Preprocessing
The data set consists of T1 scans from the multi-modal BraTS 2022 glioma segmentation challenge [4, 5, 6, 9]. The data is a retrospective collection of brain tumor MR scans originating from multiple different institutions using different equipment and imaging protocols but all working under standard clinical conditions. For each patient, the challenge organizers provide T1 scans of brains affected by tumors, masks of the tumor regions, masks of healthy tissue regions, combined masks of the tumor and healthy regions as well as "T1-voided" scans, in which the combined masks are cropped out from the original T1 image. For our approach, we use the T1 images and the healthy region masks during training. As our final goal is to learn to replace healthy tissue, for training, we crop out the healthy masks from the T1 images to obtain a masked T1 image in which only a healthy region is missing. We remove the top and bottom 0.1 percentile of voxel intensities and normalize the scans to voxel values between 0 and 1. As our model is constructed for 2D data, we slice each 3D scan into axial slices that are then cropped to a size of [224, 224]. Only background pixels are affected by the cropping. For training, we only use slices showing a non-zero mask. The provided training set consists of 1251 subjects which were used for a two-fold cross-validation, each using 1241 subjects for training and 10 for validation.
@@ 3.3 Postprocessing and Evaluation
After sampling, we apply a Gaussian filter with a standard deviation $sigma=1.075$ for the Gaussian kernel to smooth borders between the slices. In addition to the training data, the challenge organizers provided a validation set of 219 subjects that could be evaluated using our model and submitted to the Synapse [17] platform to obtain structural similarity index measure (SSIM), peak signal-to-noise ratio (PSNR) and mean squared error (MSE). For submission to the Synapse platform, each model output is normalized to the intensity range of the corresponding input image from which the first and last 0.5 percentile were clipped. After submission of the final model as a MLCube [18], it was evaluated by the challenge organizers on a non-public test set of 568 cases.
@@ 4.1 Cross-validation using Train-Validation Splits from Training Data Set
Figure 4: Axial slices of images from our validation set. The masked slice, the mask, the generated sample, and the ground truth are provided. In general, the 2D generated samples are of good quality. Moreover, they show a high similarity to the corresponding ground truth. The SSIM, MSE and PSNR that are reported refer to the comparison of the whole generated volume (obtained by stacking of the generated slices) and the corresponding ground truth volume.
Figure 4 provides examples of a given 2D masked slice, the corresponding mask, the generated sample and the ground truth. In 2D, our approach manages to generate consistent results. However, if we look at the whole 3D volume, stripe artifacts appear in the sagittal and coronal plane if multiple sequential axial slices are replaced within the volume. This is shown in Figure 5. One possibility to reduce the visibility of these artifacts is the application of a Gaussian filter, also demonstrated in Figure 5.
Figure 5: Exemplary image of our validation set in the coronal (top) and sagittal (bottom) view. We observe stripe artifacts (blue boxes) due to the stacking of multiple axial slices. These artifacts can be reduced by applying a Gaussian filter (standard deviation $sigma=1.075$ for Gaussian kernel).
@@ 4.2 Validation on Synapse Server
Table 2: Results of the evaluation of our model’s performance on the validation set, evaluated on the Synapse server.
@@
4.3 Test Set Scores
Model 2 was submitted as a MLCube for the evaluation on the non-public test set. The evaluation was performed by the challenge organizers. Table 3 provides the obtained results. Again, SSIM, PSNR and MSE are only evaluated for the inpainted region. Using this model, we achieve the second place in the inpainting challenge.
Table 3: Results of the evaluation of our model on the non-public test set.
SSIM PSNR MSE
Model 2 0.8271 [±$[\pm 0.1308]$ 20.4949 [±$[\pm 3.1141]$ 0.0115 [±$[\pm 0.0096]$
@@ 5 Discussion
We present an approach for healthy tissue inpainting based on 2D DDPMs. Taking advantage of the excellent image synthesis performance of DDPMs, we apply these models to masked brain MR images. Image context of the masked image is provided by concatenating the masked image to the input of the diffusion model in every time step $t$. These concatenated images are created to guide the denoising process. As only healthy tissue is cropped out, the ground truth is available during training. Regarding the validation, we achieve similar results in terms of SSIM, PSNR and MSE as the other challenge participants. Our model creates good results in 2D, but a drawback of our method is that the stacking of the 2D slices can lead to stripe artifacts, which we tried to reduce using a Gaussian filter. As an alternative, we tried to avoid the stripe artifacts using a sequential approach, where we included the previous ground truth slice as an additional channel that was concatenated with the noisy image during training. During sampling, the previously sampled slice was concatenated with the noisy image. However, this approach still needs modifications, as it seemed as if the errors accumulated over the sequence of slices. An advantage of the model we used for this challenge is that due to the two-dimensionality, it needs less computational resources than 3D models and shows a stable training progression. Nevertheless, in future, we aim to expand our approach to a 3D model, e.g., similar to [19]. So far we experimented with a 3D model using the same framework as presented in this paper, but it was limited regarding memory consumption, which led to limitations in network architecture tuning. A drawback of diffusion models in general is the slow sampling speed compared to other generative models. We aim to improve this using e.g., Heun sampling schemes [20] or consistency models [21]. We conclude that our method provides a baseline that shows a good performance for the task, resulting in the second place in the competition, and has potential to be fine-tuned using different architectural and sampling changes.